Fermi National Laboratory
 

Operation Summary for August 23, 2010 - August 25, 2010

 

Monday August 23

The day shift began with Operations monitoring the Tevatron (TeV) store 7999 (47.11E30) and the Recycler stash (202.1E10), with the Antiproton Source (Pbar) stacking (15.8E10), with the Linac, Booster and the Main Injector (MI) operating normally, with MiniBooNE down due to horn work, with the NuMI experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7999 at 11:30 AM and then operators prepared the CDF and D0 collision halls for access. A Linac low energy RF station (LRF2) tripped on a breaker failure at 12:02 PM. CDF and D0 came out of access at 1:40 PM. Operators began a TeV checkout that continued until the end of shift. Linac personnel completed their repairs on the LRF2 breaker at 2:26 PM. Technicians took advantage of the Linac downtime to work on several MI and NuMI kicker supplies.

During the evening shift, Operations put the TeV into HEP shot setup at 5 PM and established store 8003 at 6:15 PM with an average initial luminosity (AIL) of 183.7E30. Antiproton transfers to the Recycler were held off for a short time due to work on a mixing valve in the MI LCW system.

Tuesday August 24

The midnight shift began with Operations monitoring store 8003 (83.76E30) and the stash (218.7E10), with Pbar stacking (15.9E10), and with NuMI taking beam.

During the midnight shift, Operations terminated store 8003 at 2:17 AM, put the TeV into shot setup, and established store 8004 at 3:42 AM with an AIL of 212.36E30.

A transfer line magnet tripped off on an over-temperature indication, which caused a small amount of downtime.

During the day shift, TeV personnel completed their separator scans by 9:09 AM. Operations terminated store 8004 at 10:04 AM. Operations prepared the CDF and D0 collision halls for access; they came out of access at 1 PM. Operations conducted a TeV checkout that lasted until the end of shift. A MI RF station (MIRF18) faulted and a tech changed its modulator. NuMI personnel accessed their absorber area for a short time.

During the evening shift, Operations put the TeV into shot setup at 4:28 PM. Operations established store 8009 at 5:23 PM with an AIL of 212E30. LRF1 tripped off; operators called in a Linac technician for assistance. The tech determined that the low level RF needed to be rebooted.

Wednesday August 25

The midnight shift began with Operations monitoring store 8009 (88.89E30) and the stash (144.7E10), with Pbar stacking (31.1E10), and with MiniBooNE taking beam.

During the midnight shift, LRF5 tripped off and it took a few minutes to come back up. The Recycler Pelletron energy regulation loop got out of sync. Operators corrected it, but it took a short while to cool the beam back down. MIRF6 faulted off. After consulting with MI personnel, it was decided to leave the station off and reconfigure slipstacking to run without the station. Operations terminated store 8009 at 5:52 AM, put the TeV into shot setup at 6:05 AM, and established store 8010 at 7:09 AM with an AIL of 227.15E30.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow some MI studies and RF work today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/25/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory